# Memory Is Not the Bottleneck: Cost-Efficient Continual Learning via Weight Space Consolidation

### v3

@@ 6 Concluding Remarks
Main contributions   In many real-world settings—such as training with foundation models—storage is abundant while GPU time is costly. This challenges the long-standing assumption in continual learning that exemplar memory is the primary bottleneck. In this paper, we revisit CL under the more realistic regime of sufficient exemplar memory and show that even simple baselines like Replay become highly competitive in both performance and cost. Building on this insight, we propose a lightweight method—Weight Space Consolidation—that operates directly in weight space through selective parameter resets and running weight averaging. Our method consistently outperforms prior algorithms across class-incremental learning and LLM continual instruction tuning, while maintaining computational costs on par with Replay. Together, our findings suggest a shift in focus for future continual learning research: from optimizing memory efficiency toward maximizing computational efficiency under realistic deployment constraints.
Limitations   While our method demonstrates strong performance and cost-efficiency under sufficient exemplar memory, it assumes access to a representative subset of past data, which may not always be feasible in privacy-sensitive or streaming-only environments. Additionally, our analysis primarily focuses on class-incremental learning and continual instruction tuning with relatively clean task boundaries. Future work may explore how the proposed weight-space strategies generalize to more complex settings such as task-agnostic CL, online CL, or continual reinforcement learning.
@@ Appendix A Appendix
A.1 Sufficient Exemplar Memory Regime
Summary
In the sufficient exemplar memory regime, the replay buffer becomes an excellent proxy for the union of past tasks, which (1) mitigates forgetting (see Section A.1.1) but (2) reduces the effective learning signal for novel information (task), causing high node-reuse and low plasticity (see Section A.1.2). Section 5 empirically quantified this trade-off across different replay ratios on CIFAR-100 and ImageNet-100.
A.1.1 Stability: Sufficient exemplar memory mitigates catastrophic forgetting
@@
Similarly, with a large enough exemplar memory, the replayed data for previous tasks closely approximates their true distributions ($\tilde{P}_{j}\approx P_{j}$ for all $j<t$). Therefore training on $D_{t}\cup\mathcal{M}_{1:t-1}$ (its distribution eq. 2) yields a risk
$\tilde{R}_{1:t}(\theta)\approx\alpha_{t}\mathbb{E}_{P_{t}}\bigr{[}\ell(\theta;x)\bigr{]}+(1-\alpha_{t})\mathbb{\sum}_{j=1}^{t-1}\hat{\pi}_{j}\mathbb{E}_{\tilde{P_{j}}}\bigr{[}\ell(\theta;x)\bigr{]},$ (8)
where $\hat{\pi}_{j}$ indicates the empirical importance (i.e., size) of the tasks. Here, we may bound $|\tilde{R}_{1:t}(\theta)-R_{1:t}(\theta)|\leq\epsilon$ under standard assumptions (e.g., Lipschitz continuity in $\theta$ [34], use of representative samples during empirical risk minimization), with $\epsilon$ shrinking as the replay buffer size $K$ increases, leading to $\tilde{P_{j}}\approx P_{j}$. Assuming $R_{1:t}$ is $\mu$-strongly convex in a neighbourhood of $\theta^{*}_{1:t}$ (i.e., locally strong convex), this risk gap implies $\lVert\tilde{\theta}^{*}_{1:t}-\theta^{*}_{1:t}\rVert\leq\sqrt{2\epsilon/\mu}$ [18, 16]. Intuitively, if the two risk surfaces are proximate, their minimizers are also close in the parameter space [4, 59].
@@
the measure of how replay-based parameters after task $t$ perform on task $j$ compared to the parameters after task $j$. Here, $\Delta_{j\rightarrow t}$ remains small if $\tilde{\theta}^{*}_{1:t}\approx\theta_{1:t}^{*}$. Naturally, since $\theta^{*}_{1:t}$ is a reliable minimizer on all tasks, the small parameter drift ensures that the performance of the model trained under sufficient exemplar memory does not degrade on earlier tasks - i.e., forgetting is reduced as exemplar memory becomes larger [48, 5].
Experimentally, Figure 6 validates that increasing the exemplar memory size can reduce forgetting (hence improving stability). Here, forgetting is measured as the average over all previously learned tasks of the drop from each task’s best-ever accuracy to its accuracy after the final task [78]. Formally:
@@
A.1.2 Plasticity: Sufficient memory deteriorates model’s capacity to learn new tasks
@@
We claim that while sufficient exemplar memory improves stability, it inevitably suppresses plasticity [13]. Let $\theta^{(t-1)}$ and $\theta^{(t)}$ denote the network parameters right before and right after training on task $t$. Denote by $g_{t}(x,y)=\nabla_{\theta}\ell(\theta^{(t-1)};x,y)$ the per–example gradient when we begin task $t$. Under the hybrid distribution of equation 2, the expected update direction is
$\underbrace{\mathbb{E}_{x\sim P_{\text{train}}^{(t)}}[g_{t}(x)]}_{\triangleq\;\bar{g}_{t}}\;=\;\alpha_{t}\,\bar{g}_{t}^{(\text{new})}\;+\;(1-\alpha_{t})\,\bar{g}_{t}^{(\text{past})},$ (10)
@@
using the cosine similarity [14, 38]. When the buffer is sufficient, $P^{(t)}_{\text{past}}$ (i.e., past task data at task step $t$ stored in the replay memory) and $P^{(t-1)}_{\text{train}}$ (i.e., the mixed train data at task step $t-1$) are close, and by definition the distribution $P^{(t)}_{\text{train}}$ at task $t$ is similar to the distribution $P^{(t-1)}_{\text{train}}$ at task $t-1$; hence $\rho_{t}\!\approx\!1$. Consequently, the effective step taken during task $t$
$\|\theta^{(t)}-\theta^{(t-1)}\|\;=\;\eta\,\|\bar{g}_{t}\|\;\leq\;\eta\bigl{[}\alpha_{t}+(1-\alpha_{t})\bigr{]}\|\bar{g}_{t}^{(\text{new})}\|\;=\;\eta\,\|\bar{g}_{t}^{(\text{new})}\|,$ (12)
where $\eta$ is the learning rate. Here, Equation 12 shrinks monotonically with $\alpha_{t}$ because $\bar{g}_{t}^{(\text{new})}$ and $\bar{g}_{t}^{(\text{past})}$ are almost colinear. In the limit $\alpha_{t}\!\to\!0$ (i.e. the current data is overwhelmed by replay) the update direction collapses onto the span of previous gradients [74, 33], yielding
[…]
which encourages the network to remain dormant (i.e., loss of plasticity [13]) between tasks.
This analysis connects with the node-reuse phenomenon studied in [39], which claimed that when two sequential tasks are similar–as in the case of $P^{t}_{\text{train}}$ and $P^{(t-1)}_{\text{train}}$ under the sufficient memory regime–models tend to reuse their nodes (i.e., loss of plasticity), and also aligns with the model’s underperformance in training under extreme replay (i.e., full data is provided as exemplar memory) as reported in [55].
The results of Figure 7 depict how a model’s plasticity, more specifically its ability to learn new tasks, diminishes as exemplar memory becomes sufficient. Here, the plasticity was measured using the model’s average accuracy on new tasks. In Section 3.2, we also provided another analysis focused on the train loss, where we observe that under sufficient exemplar memory, convergence becomes difficult. In the sequel, we investigate this phenomenon more deeply.
A.1.3 Discussion: New challenges under sufficient exemplar memory
@@
This combined dataset introduces several side effects. First, as discussed in Section 3.2, stability increases, hence reducing forgetting (Section A.1.1). On the other hand, plasticity worsens (Section A.1.2). In this section, we investigate whether the sufficient memory regime poses additional issues.
@@
Conventional continual learning methods typically address forgetting by assuming that new data is markedly different from prior data and focusing on preserving performance on previously learned tasks. However, in the sufficient exemplar memory regime, where all tasks are presented simultaneously within a combined dataset, these methods fall short. They are not equipped to handle the multi-task learning dynamics and the associated balancing issues that emerge when the model must integrate and harmonize learning signals from a diverse set of tasks. We believe a more thorough investigation is required, and we set this as a key objective of our future work.
@@
In this section, we elaborate on our method weight space consolidation, highlighting the role of each component. The proposed method is a combination of two weight-space operations. (1) rank-based parameter reset [68] and (2) weight averaging [66].
Rank-based weight reset
@@
The weight reset step selects and resets redundant parameters in adapting to the new task. The aim of this procedure is to find the minimal set of parameters that are capable of adapting to the new $t^{th}$ task, and reinitializing the redundant parameters to the mixed value of the previous task model $\theta_{prev}$ and the current model $\theta$ value using Equation 5, which helps recover learned features. In this process, we use a simple metric Equation 4 that uses the parameter’s first and second moments to measure its importance [35, 3, 51, 26]. In Section A.3, we also compare alternative measures (e.g., parameter drift) and discuss their limitations.
@@
Using this metric, we retain the top-$Q\%$ parameters that have largely drifted during the current task and reset the dormant parameters using eq. 5. The idea of resetting model weights is not novel in the continual learning literature, but most have focused on improving plasticity [13]. Conversely, we improve both plasticity and stability using a mixing technique (eq. 5, integrating previous and present tasks. In section A.4, we provide experimental results on the effect of $Q$, the percentage of weight reset on model performance (see fig. 5). Furthermore, we study alternative variations of our reset method, namely (1) different parameter importance metrics (see Table 4), (2) different reset methods and reset frequencies (see Table 4). We find that resetting more frequently (e.g., per epoch) displays significant gains in performance especially under limited memory settings. We discuss this observation on Section A.4.
Weight Average
@@
The weight-averaging step aggregates various learned signals into a single model, allowing for faster convergence and improved generalization [28]. The underlying idea is that the model weights can address catastrophic forgetting, functioning as a substitute for replay memory. Using this, we aim to fill the gap between the full exemplar memory setting (i.e., joint-training) and the sufficient exemplar memory setting. Another motive behind this approach is the new challenges (see section 3.1) rooted in training with a mixture of datasets $D_{t}\cup\mathcal{M}_{1:t-1}$, which generally requires more training steps for convergence. In Section A.4, we empirically observe that the sufficient exemplar memory setting complicates training, requiring more training epochs for convergence (see Figure 2(b)). The weight averaging technique is also an emerging practice in the continual learning literature [46, 37]. However, such works merge the model weights after training (i.e., offline merging [15]), which requires the storage of multiple model weights (proportional to the number of tasks). On the other hand, our method uses a moving average model that is updated during training (i.e., online merging).
@@ A.3 Measuring Parameter Importance
In our work, we measure a parameter’s contribution to learning a new task by using a moment-based score (eq. 4). However, there are several alternative approaches we could take. In this section, we investigate the alternative measures that could be used to measure a parameter’s importance to learning a new task.
@@
which captures how strongly the log-likelihood reacts to perturbations of $\theta[l]$. This idea underpins many existing continual learning methods (e.g., EWC [36], MAS [1]). However, Fisher scores are usually difficult to compute real-time and lack scalability.
Another metric we can use is the Hessian estimate [73, 12]. Second-order methods replace $F_{l}$ by a Hessian diagonal estimate, which we can efficiently obtain with Hutchinson’s trick [25]:
@@
where each Hutchinson probe costs a single Hessian–vector product. The absolute curvature $s_{l}=\lvert\widehat{h}_{l}\rvert$ can then serve as our importance score. However, similar to the Fisher matrix, Hessians are exceptionally costly to compute, especially for large models. In Section 5, we investigate the efficacy of these metrics as a parameter importance measure.
Inter-task and Intra-task
@@
Inter-task Parameter Drift
@@
where $\theta_{t}[l]$ denotes the $l^{th}$ parameter after training on the current task $t$, and $\theta_{t-1}[l]$ represents the $l^{th}$ parameter after training on the previous task $t-1$. A large value indicates that the parameter is highly task-specific, while a small value suggests robustness across tasks.
Intra-task Parameter Drift

### v5

@@ 6 Concluding Remarks
This paper challenges the long-standing assumption in continual learning (CL) that exemplar memory is the primary bottleneck, arguing that in modern deployments, GPU cost is the true constraint. We investigate the consequences of having sufficient memory—a practical regime where forgetting is largely solved but full retraining from scratch remains prohibitively expensive—and demonstrate that the central challenge shifts from mitigating forgetting to overcoming a loss of plasticity. To address this, we propose Weight Space Consolidation, a lightweight method combining parameter resets and weight averaging to navigate this new stability-plasticity trade-off. Our method is empirically validated across image classification and LLM instruction tuning, where it outperforms strong baselines at a fraction of the cost, establishing a scalable and cost-efficient alternative to full retraining. Ultimately, our findings call for a crucial shift in focus for future CL research: from optimizing for unrealistic memory constraints toward designing computationally efficient algorithms for the real-world scenarios of today.
Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) with a grant funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
Ethics Statement
The authors acknowledge and concur with the ICLR Code of Ethics, namely in its pursuit of (1) human well-being, (2) high standards of scientific excellence, (3) consideration for the societal impacts (i.e., harms) of AI, (4) honesty & trustworthiness, (5) fairness, (6) mutual respect for other researchers’ works, (7) privacy, and (8) confidentiality.
Reproducibility Statement
For reproducibility, we provide the source code, experimental guidelines, and the scripts used in our experiments. Please refer to the README.md file in the supplementary materials on how to reproduce our experiments. We also used a fixed seed setting, which is implemented in the source code. We also include notebook (.ipynb) files to reproduce the figures appearing in our paper. Lastly, in Section˜5.1 and Section˜A.5, we thoroughly explain how our method and its experiments are implemented.
@@ Appendix A Appendix
A.1 abundant exemplar memory Regime
Summary.
In the abundant exemplar memory regime, the replay buffer becomes an excellent proxy for the union of past tasks, which (1) mitigates forgetting (see Section˜A.1.1) but (2) reduces the effective learning signal for novel information (task), causing high node‑reuse and low plasticity (see Section˜A.1.2). Section˜5 empirically quantified this trade‑off across different replay ratios on CIFAR‑100 and ImageNet‑100.
A.1.1 Stability: abundant exemplar memory mitigates catastrophic forgetting
@@
Similarly, with a large enough exemplar memory, the replayed data for previous tasks closely approximates their true distributions ($\tilde{P}_{j}\approx P_{j}$ for all $j<t$). Therefore training on $D_{t}\cup\mathcal{M}_{1:t-1}$ (its distribution eq.˜2) yields a risk
$\tilde{R}_{1:t}(\theta)\approx\lambda\mathbb{E}_{P_{t}}\bigr[\ell(\theta;x)\bigr]+(1-\lambda)\mathbb{\sum}_{j=1}^{t-1}\hat{\pi}_{j}\mathbb{E}_{\tilde{P_{j}}}\bigr[\ell(\theta;x)\bigr],$ (8)
where $\hat{\pi}_{j}$ indicates the empirical importance (i.e., size) of the tasks. Here, we may bound $|\tilde{R}_{1:t}(\theta)-R_{1:t}(\theta)|\leq\epsilon$ under standard assumptions (e.g., Lipschitz continuity in $\theta$ (Khromov and Singh, 2024), use of representative samples during empirical risk minimization), with $\epsilon$ shrinking as the replay buffer size $K$ increases, leading to $\tilde{P_{j}}\approx P_{j}$. Assuming $R_{1:t}$ is $\mu$‑strongly convex in a neighbourhood of $\theta^{*}_{1:t}$ (i.e., locally strong convex), this risk gap implies $\lVert\tilde{\theta}^{*}_{1:t}-\theta^{*}_{1:t}\rVert\leq\sqrt{2\epsilon/\mu}$ (Fornasier et al., 2021; Escande, 2024). Intuitively, if the two risk surfaces are proximate, their minimizers are also close in the parameter space (Beer et al., 1992; van de Geer and Wainwright, 2017).
@@
the measure of how replay-based parameters after task $t$ perform on task $j$ compared to the parameters after task $j$. Here, $\Delta_{j\rightarrow t}$ remains small if $\tilde{\theta}^{*}_{1:t}\approx\theta_{1:t}^{*}$. Naturally, since $\theta^{*}_{1:t}$ is a reliable minimizer on all tasks, the small parameter drift ensures that the performance of the model trained under abundant exemplar memory does not degrade on earlier tasks - i.e., forgetting is reduced as exemplar memory becomes larger (Merlin et al., 2022; Brignac et al., 2023).
Experimentally, Figure˜4 validates that increasing the exemplar memory size can reduce forgetting (hence improving stability). Here, forgetting is measured as the average over all previously learned tasks of the drop from each task’s best‐ever accuracy to its accuracy after the final task (Zhou et al., 2023a). Formally:
@@
A.1.2 Plasticity: abundant exemplar memory deteriorates model’s capacity to learn new tasks
@@
$\lambda$
We claim that while abundant exemplar memory improves stability, it inevitably suppresses plasticity (Dohare et al., 2024). Let $\theta^{(t-1)}$ and $\theta^{(t)}$ denote the network parameters right before and right after training on task $t$. Denote by $g_{t}(x,y)=\nabla_{\theta}\ell(\theta^{(t-1)};x,y)$ the per–example gradient when we begin task $t$. Under the hybrid distribution of equation 2, the expected update direction is
$\underbrace{\mathbb{E}_{x\sim P_{\text{train}}^{(t)}}[g_{t}(x)]}_{\triangleq\;\bar{g}_{t}}\;=\;\lambda\,\bar{g}_{t}^{(\text{new})}\;+\;(1-\lambda)\,\bar{g}_{t}^{(\text{past})},$ (10)
@@
using the cosine similarity (Du et al., 2018; Lee et al., 2021). When the buffer is abundant, $P^{(t)}_{\text{past}}$ (i.e., past task data at task step $t$ stored in the replay memory) and $P^{(t-1)}_{\text{train}}$ (i.e., the mixed train data at task step t1t-1) are close, and by definition the distribution $P^{(t)}_{\text{train}}$ at task $t$ is similar to the distribution $P^{(t-1)}_{\text{train}}$ at task t1t-1; hence $\rho_{t}\!\approx\!1$. Consequently, the effective step taken during task $t$
$\|\theta^{(t)}-\theta^{(t-1)}\|\;=\;\eta\,\|\bar{g}_{t}\|\;\leq\;\eta\bigl[\lambda+(1-\lambda)\bigr]\|\bar{g}_{t}^{(\text{new})}\|\;=\;\eta\,\|\bar{g}_{t}^{(\text{new})}\|,$ (12)
where $\eta$ is the learning rate. Here, Equation˜12 shrinks monotonically with $\lambda$ because $\bar{g}_{t}^{(\text{new})}$ and $\bar{g}_{t}^{(\text{past})}$ are almost colinear. In the limit $\lambda\!\to\!0$ (i.e. the current data is overwhelmed by replay) the update direction collapses onto the span of previous gradients (Yu et al., 2020; Kendall et al., 2018), yielding
$\|\theta^{(t)}-\theta^{(t-1)}\|\;\xrightarrow[\lambda\to 0]{}\;0,$ (13)
which encourages the network to remain dormant (i.e., loss of plasticity (Dohare et al., 2024)) between tasks.
This analysis connects with the node‑reuse phenomenon studied in (Lee et al., 2022), which claimed that when two sequential tasks are similar–as in the case of $P^{t}_{\text{train}}$ and $P^{(t-1)}_{\text{train}}$ under the abundant exemplar memory regime–models tend to reuse their nodes (i.e., loss of plasticity), and also aligns with the model’s underperformance in training under extreme replay (i.e., full data is provided as exemplar memory) as reported in (Rolnick et al., 2019).
The results of Figure˜5 depict how a model’s plasticity, more specifically its ability to learn new tasks, diminishes as exemplar memory becomes abundant. Here, the plasticity was measured using the model’s average accuracy on new tasks. In Section˜3.2, we also provided another analysis focused on the train loss, where we observe that under abundant exemplar memory, convergence becomes difficult. In the sequel, we investigate this phenomenon more deeply.
A.1.3 Discussion: New challenges under abundant exemplar memory
@@
This combined dataset introduces several side effects. First, as discussed in Section˜3.2, stability increases, hence reducing forgetting (Section˜A.1.1). On the other hand, plasticity worsens (Section˜A.1.2). In this section, we investigate whether the abundant exemplar memory regime poses additional issues.
@@
Conventional continual learning methods typically address forgetting by assuming that new data is markedly different from prior data and focusing on preserving performance on previously learned tasks. However, in the abundant exemplar memory regime, where all tasks are presented simultaneously within a combined dataset, these methods fall short. They are not equipped to handle the multi-task learning dynamics and the associated balancing issues that emerge when the model must integrate and harmonize learning signals from a diverse set of tasks. We believe a more thorough investigation is required, and we set this as a key objective of our future work.
@@
In this section, we elaborate on our method weight space consolidation, highlighting the role of each component. The proposed method is a combination of two weight-space operations. (1) ranking-based parameter reset (Yadav et al., 2024) and (2) weight averaging (Wortsman et al., 2022).
@@
Ranking-based weight reset.
@@
The weight reset step selects and resets redundant parameters in adapting to the new task. The aim of this procedure is to find the minimal set of parameters that are capable of adapting to the new $t^{th}$ task, and reinitializing the redundant parameters to the mixed value of the previous task model $\theta_{prev}$ and the current model $\theta$ value using Equation˜5, which helps recover learned features. In this process, we use a simple metric Equation˜4 that uses the parameter’s first and second moments to measure its importance (Kingma and Ba, 2017; Balles and Hennig, 2020; Molchanov et al., 2019; Hwang, 2024). In Section˜A.3, we also compare alternative measures (e.g., parameter drift) and discuss their limitations.
@@
Using this metric, we retain the top-$Q\%$ parameters that have largely drifted during the current task and reset the dormant parameters using eq.˜5. The idea of resetting model weights is not novel in the continual learning literature, but most have focused on improving plasticity (Dohare et al., 2024). Conversely, we improve both plasticity and stability using a mixing technique (eq.˜5, integrating previous and present tasks. In section˜A.4, we provide experimental results on the effect of $Q$, the percentage of weight reset on model performance (see fig.˜11). Furthermore, we study alternative variations of our reset method, namely (1) different parameter importance metrics (see Table˜4), (2) different reset methods and reset frequencies (see Table˜5). We find that resetting more frequently (e.g., per epoch) displays significant gains in performance especially under limited memory settings. We discuss this observation on Section˜A.4.
Weight Average.
@@
The weight-averaging step aggregates various learned signals into a single model, allowing for faster convergence and improved generalization (Izmailov et al., 2018). The underlying idea is that the model weights can address catastrophic forgetting, functioning as a substitute for replay memory. Using this, we aim to fill the gap between the full exemplar memory setting (i.e., joint-training) and the abundant exemplar memory setting. Another motive behind this approach is the new challenges (see section˜3.1) rooted in training with a mixture of datasets $D_{t}\cup\mathcal{M}_{1:t-1}$, which generally requires more training steps for convergence. In Section˜A.4, we empirically observe that the abundant exemplar memory setting complicates training, requiring more training epochs for convergence (see Figure˜2(b)). The weight averaging technique is also an emerging practice in the continual learning literature (Marouf et al., 2025; Kozal et al., 2024). However, such works merge the model weights after training (i.e., offline merging (Dziadzio et al., 2024)), which requires the storage of multiple model weights (proportional to the number of tasks). On the other hand, our method uses a moving average model that is updated during training (i.e., online merging).
Why does Weight Space Consolidation work?
Our starting point is the abundant memory regime, where abundant exemplars reduce forgetting but drive the optimizer toward sharp, over-specialized minima that harm plasticity. Weight Space Consolidation is designed to counteract this effect with two complementary weight-space operations. First, targeted re-initialization of dormant parameters restores unused capacity, enabling the network to escape locally saturated directions and learn new tasks. Second, in-situ weight averaging biases training toward flatter regions of the loss landscape, a mechanism that prior work has linked to improved generalization and reduced forgetting Izmailov et al. (2018); Cha et al. (2021; 2020). Our experiments further show that this averaging step induces sparsity in the effective parameter usage, and sparsity is known to enhance plasticity by preventing the model from over-relying on a small set of critical parameters Golkar et al. (2019); Dohare et al. (2024). Together, these effects provide an explanation of why our proposed method can simultaneously recover plasticity and maintain stability in the abundant-memory regime.
@@ A.3 Measuring Parameter Importance
In our work, we measure a parameter’s contribution to learning a new task by using a moment-based score (eq.˜4). However, there are several alternative approaches we could take. In this section, we investigate the alternative measures that could be used to measure a parameter’s importance to learning a new task.
@@
which captures how strongly the log‑likelihood reacts to perturbations of $\theta[l]$. This idea underpins many existing continual learning methods (e.g., EWC (Kirkpatrick et al., 2017), MAS (Aljundi et al., 2018)). However, Fisher scores are usually difficult to compute real-time and lack scalability.
Another metric we can use is the Hessian estimate (Yu et al., 2021; Chong et al., 2023). Second‑order methods replace $F_{l}$ by a Hessian diagonal estimate, which we can efficiently obtain with Hutchinson’s trick (Hutchinson, 1986):
@@
where each Hutchinson probe costs a single Hessian–vector product. The absolute curvature $s_{l}=\lvert\widehat{h}_{l}\rvert$ can then serve as our importance score. However, similar to the Fisher matrix, Hessians are exceptionally costly to compute, especially for large models. In Section˜5, we investigate the efficacy of these metrics as a parameter importance measure.
@@
Inter-task Parameter Drift.
@@
where $\theta_{t}[l]$ denotes the $l^{th}$ parameter after training on the current task $t$, and $\theta_{t-1}[l]$ represents the $l^{th}$ parameter after training on the previous task t1t-1. A large value indicates that the parameter is highly task-specific, while a small value suggests robustness across tasks.
Intra-task Parameter Drift.